OpenStudy (anonymous):

construct a polynomial of degree 4 that touches the x axis at -3 and stays on the positive side of the axis and also goes through the x axis at 2 and 5. Also determine the constant multiplier that makes the function have (1,192) on the graph.

OpenStudy (amistre64):

what will be your 4 roots, since a degree 4 needs 4 roots

OpenStudy (anonymous):

how do I figure that out?

OpenStudy (amistre64):

you need to see where it ineracts with the x axis

OpenStudy (anonymous):

that would be -3, 2 and 5 ?

OpenStudy (anonymous):

but Im not sure about the 4th

OpenStudy (amistre64):

yes, but with a little modification think of "just touching" as going into the axis at -3, and out of the axis at -3, giving us a multiple root at -3 roots: -3,-3,2,5

OpenStudy (anonymous):

oh ok. that makes sense

OpenStudy (amistre64):

now make a product with your roots, by subtracting "x" from each root. tell me if that makes any sense

OpenStudy (anonymous):

-3x^4-3x^3+2x^2+5x

OpenStudy (amistre64):

well, that might be correct, but its a step or 2 away

OpenStudy (amistre64):

does the product of the roots make sense?

OpenStudy (amistre64):

(-3-x) (-3-x) (2-x) (5-x)

OpenStudy (anonymous):

I guess Im not sure what you mean

OpenStudy (anonymous):

do the products need to equal something thats the same?

OpenStudy (amistre64):

the "multiply by zero" property makes y=0 whenever we use a root. 2 (-3-x) (-3-x) (2-x) (5-x) = 0 0 -3 (-3-x) (-3-x) (2-x) (5-x) = 0 0 -3 (-3-x) (-3-x) (2-x) (5-x) = 0 0 5 (-3-x) (-3-x) (2-x) (5-x) = 0 0 by subtracting x from each root, and multiplying the terms together, we get a polymonial

OpenStudy (amistre64):

y = (-3-x) (-3-x) (2-x) (5-x) y = (x^2-6x+9) (x^2-7x +10) y = x^4 - x^3 - 23x^2 - 3x + 90 ; this will be your basic polynomial with the given roots but we have to make one more modification ...

OpenStudy (amistre64):

we do not want to "add" anything to the basic structure, otherwise we alter the root: 0+b = b. and not 0 so we can only multiply this set up to modify it to meet our needs

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

what do we get when x = 1 in our basic setup? y = 1 - 1 - 23 - 3 + 90 = 64 since we want x=1, y=192, and we can only multiply to modify this we have 64n = 192 n = 192/64 = 3 so lets multiply the basic setup by 3: y = 3x^4 - 3x^3 - 69x^2 - 9x + 270 is the poly we desire

OpenStudy (anonymous):

its actually 1192

OpenStudy (amistre64):

lol, same process, different "n"

OpenStudy (amistre64):

the point in question is: (1, 192) ? or, (1, 1192)?

OpenStudy (anonymous):

you're right 1, 192

OpenStudy (amistre64):

yay!! i can stave of senility for another day :)

OpenStudy (anonymous):

lol

OpenStudy (amistre64):

a textbook method is similar to mine, except they propose you subtract the root from x to make the product (x--3)(x--3)(x-2)(x-5) (x+3)(x+3)(x-2)(x-5) ... gets you the same answer, but i find the subtract a negative and switch to a plus a little useless .... just an added step is all

OpenStudy (anonymous):

ok cool thank you

OpenStudy (amistre64):

good luck ;)

OpenStudy (anonymous):

do you know how I would determine the coordinates of the turning points?

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

you know one of the turning points already (-3,0) has to be a turning point since it only touches the line and doesnt go thru it.

OpenStudy (amistre64):

are you familiar with derivatives? from calculus?